### v1

# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

Abstract.
High-Level Synthesis enables the rapid prototyping of hardware accelerators, by combining a high-level description of the functional behavior of a kernel with a set of micro-architecture optimizations as inputs. Such pragmas may describe the pipelining and replication of units, or even higher level transformations for HLS such as automatic data caching using the AMD/Xilinx Merlin compiler. Selecting the best combination of pragmas, even within a restricted set, remains particularly challenging and the typical state-of-practice uses design-space exploration to navigate this space. But due to the highly irregular performance distribution of pragma configurations, typical DSE approaches are either extremely time consuming, or operating on a severely restricted search space.
In this work we propose a framework to automatically insert HLS pragmas in regular loop-based programs, supporting pipelining, unit replication (coarse- and fine-grain), and data caching. We develop a simple analytical performance and resource model as a function of the input program properties and pragmas inserted. We prove this model provides a lower bound on the actual performance for any possible configuration. We then encode this model as a Non-Linear Program, by making the pragma configuration as unknowns of the system, which is computed optimally by solving this NLP. This approach can also be used during DSE, to quickly prune points with a (possibly partial) pragma configuration, employing this latency lower bound property. We extensively evaluate our end-to-end fully implemented system, showing it can effectively manipulate spaces of billions of designs in seconds to minutes for the kernels evaluated.
HLS, FPGA, Non-Linear Programming, Program Optimization
copyright: noneconference: Proceedings of the 2024 ACM/SIGDA International Symposium on Field Programmable Gate Arrays; March 3 –- 5, 2024 ; Monterey, CAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
High-level synthesis (HLS) (Cong et al., 2011; Zhang et al., 2008) compilers (Xilinx, 2023b; Intel, 2023; Siemens, 2023; Microchip, 2023) and source-to-source compiler for HLS (Xilinx, 2023a; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020) can reduce development time while delivering a good performance for the designs. However, achieving a satisfactory Quality of Results (QoR) often requires design-space exploration (DSE). This is because the design space, including which pragmas to insert and where, can not only contain millions of points, but typically does not present characteristics suitable for fast analytical exploration, such as convexity and regularity. Although the existing DSE methods (Sohrabizadeh et al., 2021, 2022; Wu et al., 2023) can find designs with a good QoR, it comes at a high computation cost: for example hundreds, of designs may be concretely instantiated using HLS to compute its estimated QoR during exploration (Sohrabizadeh et al., 2021).
Our main objective is to provide a system to automatically insert a set of hardware pragmas for HLS that delivers a good QoR but significantly reduces the search time needed to obtain the final design. To address this challenge, we propose NLP-DSE, a framework built on top of the AMD/Xilinx Merlin source-to-source compiler (Xilinx, 2023a). This framework automatically inserts pragmas for unrolling/parallelization, pipelining, tiling and data caching for linear programs prior to HLS. These Merlin pragmas can also be inserted using a DSE approach, such as AutoDSE (Sohrabizadeh et al., 2021), which we utilize as reference for our evaluations. However, in contrast to AutoDSE (Sohrabizadeh et al., 2021), we develop a hybrid analytical approach combined with a lightweight DSE, reducing the number of designs explored while preserving, and often even improving, the final QoR of designs produced.
To this end, we create a novel Non-Linear Programming approach to automatically insert pragmas in an existing program. We develop an analytical model combining latency and resources, targeting regular loop-based kernels (PolyBench, [n. d.]), that is parameterized by the pragma configuration. We can then designate the pragma configuration as the unknowns of this model, solving it by NLP to obtain the set of pragmas that minimizes latency. An important design principle of our approach is to ensure that the latency computed is a lower bound on the achievable latency for a given pragma configuration. This enables efficient pruning during DSE: any design predicted to have a latency lower bound higher than the best latency obtained through exploration so far is necessarily slower and does not need exploration. To overcome the fact that optimizing compilers, and the overall HLS toolchain underneath, may not apply optimizations as expected (e.g., due to insufficient resources, or limitations of the compiler’s implementation), we develop a lightweight NLP-based DSE approach, exploring parts of the design space with different types and amounts of hardware parallelisms and array partitioning factors. We make the following contributions:
  • We produce an analytical performance and resource model specifically built for AMD/Xilinx Vitis and Merlin compilers, which is amenable to optimization via non-linear programming.
  • We prove our model is a lower bound on the final latency of the design, under reasonable hypothesis. This enables fast pruning of the design space, ensuring pruned designs have higher latency.
  • We develop a NLP-based DSE approach exploiting this model, targeting regular loop-based kernels, which can significantly outperform DSE-based search approaches such as AutoDSE, delivering equal or better QoR in a significantly less search time.
  • We implement NLP-DSE as an end-to-end fully automated system and use it to conduct extensive evaluation on 47 benchmarks including kernels from linear algebra, image processing, physics simulation, graph analytics, datamining, etc. (PolyBench, [n. d.]). Our results show the ability of our approach to find in most cases a better QoR than AutoDSE, in significantly less time.
The paper is organized as follows. Section 2 motivates our approach and solution proposed. Section 3 presents our analytical performance and resource model. Section 4 delves into proving it is a lower bound on the final QoR. In Section 5, we introduce a non-linear formulation based on this model to automatically find pragma configurations by NLP optimization. Section 6 presents our lightweight DSE approach. Finally, Sections 7 and 9 are devoted to evaluating our method validating the effectiveness of our approach and presenting related work, before concluding.
2. Background and Motivation
The main objective of this work is to automatically optimize FPGA designs using HLS (Cong et al., 2011; Zhang et al., 2008). HLS has made FPGA usage more accessible, and many projects are looking to further democratize this field by automating optimizations (Sohrabizadeh et al., 2021; Lai et al., 2019; Huang et al., 2021; Ye et al., 2022; Ferretti et al., 2018) such as the AMD/Xilinx Merlin Compiler ((Cong et al., 2016b, a; Xilinx, 2023a)).
In particular, Merlin was developed to improve the performance and reduce the development time of HLS-based designs. To achieve this goal, it automatically generates data transfers between off-chip and on-chip memory and inserts important Vitis pragmas, such as for array partitioning. Merlin includes hardware directives pragmas such as ACCEL parallel <factor=x>, which creates x parallel instances, and Merlin restructures the code accordingly if the loop nest has more than x iterations. This pragma can be used for fine-grained and coarse-grained parallelization. There are also the pragmas ACCEL pipeline <II=y> for pipelining, and ACCEL tile <factor=z> for strip-mining a loop by z, enabling Merlin to insert other pragmas such as data caching in a loop with smaller trip count, matching the on-chip resources available and reducing off-chip communications. Merlin’s ACCEL cache <array=a> directive transfers all required elements of array $a$ from off-chip to on-chip to perform computations within the specified sub-region.
2.1. DSE for Pragma Insertion
Tools such as AutoDSE (Sohrabizadeh et al., 2021) and GNN-DSE (Sohrabizadeh et al., 2022) rely on Merlin and perform DSE to find which pragmas to insert and where.
Although Merlin further raises the level of abstraction and easiness to implement complex transformations such as data caching and various types of hardware replication and hence reduces the space of configurations, the space of pragmas and their parameter values is still large: as shown later in Sec. 7, this space quickly reaches billions of feasible designs even for kernels containing only a handful of loops and statements.
Furthermore, to avoid accuracy limitations of analytical performance models, DSE approaches exploring such space typically rely on partially generating the design, at least via HLS, and utilizing the QoR report as the performance estimate for a design. HLS time for highly optimized designs combining various Merlin pragmas (e.g., parallelism and caching) can quickly reach several hours per design, making the search process particularly time-consuming. One may restrict the space of pragmas considered, and especially their parameter range, to reduce search time (Sohrabizadeh et al., 2021). However this comes at the expense of the potentially improved performance that can be reached by considering the complete parameter space.
Yet such DSE remains the state-of-practice approach to deliver high-quality designs, we illustrate its performance merits with three sample benchmarks from linear algebra: GEMM, the dense generalized matrix-multiply, 2mm, which computes the matrix product $D=A*B*C$, and Gramschmidt, which computes the Gram-Schmidt process for QR decomposition. Table 1 below displays the performance (in GigaFlop/s, GF/s) of the original program from PolyBench/C using the Medium dataset size and Large for Gramschmidt, without any pragmas, and optimized by Merlin and the final performance achieved by running AutoDSE to insert Merlin pragmas, given a time budget of 20 hours and a timeout of 3 hours per HLS run. DSE time is reported in minutes.
2mm Gemm Gramsch.
GF/s Time GF/s Time GF/s Time
Original 0.10 N/A 0.07 N/A 0.14 N/A
AutoDSE 0.41 1,870 68.91 1,345 0.95 819
NLP-DSE 117.48 70 105.18 185 2.34 420
Table 1. Result of AutoDSE
This framework utilizes the HLS compilers as a black box used to select configurations that minimize the objective function. The tool is agnostic of the input program shape and if it detects that Merlin did not apply the pragmas as expected it allows the DSE to prune the design after Merlin has generated the HLS-C code.
By analyzing the space explored by the DSE for these three examples, valuable hints can be observed to improve the method. For 2mm, the fastest design found by AutoDSE is mainly optimized for a single loop body. When AutoDSE tries to optimize the second loop body, it favors the unroll factors to the power of two for the innermost loop and goes directly to the outermost loop. This does not improve performance or create configurations which are pruned. The fact that it does not try the other unrolls factors for the innermost loop before optimizing the other pragmas creates a loss of performance. For Gemm Medium and Gramschmidt Large, the DSE makes it possible to find designs with a good QoR. However the DSE wastes much time exploring too large unroll factors, which generates ponderously long synthesis times without giving any result as the HLS timeout is reached. The time spent increasing the unroll factor for certain pragmas without result does not allow the unroll factor of other pragmas to be increased, which results in missed performance.
2.2. Overview of NLP-DSE
Our primary objective is to efficiently achieve high-quality designs within a limited compilation time budget. To accomplish this we develop a DSE based on a cost model which can find the theoretical optimal configuration in a few seconds/minutes, thanks to its implementation as a Non-Linear Programming (NLP) problem which is then solved using BARON (Tawarmalani and Sahinidis, 2005). Note, however, the in-practice configuration that achieves the best performance may be different: our NLP is unable to model the complexity of the downstream toolchain and issues of resource usage/sharing, actual implementation of the pragmas by Merlin and Vitis, and simple performance bugs in these optimizing compilers may imply a theoretically best design is not implemented as such, making a design space exploration approach still required to achieve a good performance. We have therefore set up a DSE which allows the framework to explore different spaces of parallelism as well as different configuration spaces by constraining the level of parallelism. Our model is presented in Sec. 3, and we prove it is a performance lower bound in Sec. 4, an important feature to be able to prune designs during the search without the risk of losing performance. The associated NLP formulation is given in Sec. 5.
To make our approach feasible, and in particular maintain a sufficient level of accuracy in the analytical models, we target regular, loop-based computations, that is affine programs (Feautrier, 1988). Such programs have a static control-flow, that is every operation executed, and a branch taken at run-time, can be exactly modeled at the compile-time using polyhedral structures (Feautrier, 1992). This restriction improves accuracy of loop-based analytical model by providing them with exact information about the loop trip counts and dependencies.
It is important to note that our DSE process, presented in Sec. 6, fundamentally differs from AutoDSE. We first evaluate configurations that have the lowest theoretical latency, and hence typically a high level of parallelism. However AutoDSE starts from a configuration with no pragmas, which is typically feasible in terms of resource usage, and progressively increases the pragmas until resources are exceeded. Instead our NLP-DSE begins with a design featuring the maximum achievable parallelism, and then gradually reduces the level of parallelism. We illustrate this with Table 1 with the final performance and time-to-solution achieved by NLP-DSE.
As presented in Sec. 7, compared to AutoDSE, our framework locates designs with similar or higher throughput for 46 out of 47 benchmarks evaluated, with a performance improvement of 5.69x and 17.24x on average in terms of the DSE time and design throughput respectively. Only 1 design having a 7% lower throughput. In all 47 benchmarks, the time-to-solution of NLP-DSE is significantly less than AutoDSE, up to 30x faster.
3. Modeling Programs and Their Pragmas
We now present our analytical performance model. We assume the input programs are polyhedral programs (Feautrier, 1992; Girbal et al., 2006), and therefore exact loop trip counts can be computed by static analysis, similarly for all data dependencies.
3.1. Program Representation
We represent programs using a summary of their Abstract syntax tree (AST), with sufficient information to estimate latency and resource consumption by analytical modeling. Intuitively, we can build a constructor-style description of the summary AST, and then directly instantiate the complete formula for estimating e.g., latency, based on loop properties. We first introduce this representation before proving how to compute a latency lower bound with it.
We employ the code below as a running example with the pragma above the loops as AMD/Xilinx Merlin and OpenMP. For presentation simplicity, we assume each loop iterator in the program region has been renamed to a unique name, so that we can uniquely identify loops in the code by their iterator name.
1<some-pragmas-for-loop-i>
2for (i = lbi; i <= ubi; i++) {
3 <some-pragmas-for-loop-j1>
4 for (j1 = lbj1(i); j1 <= ubj1(i); j1++) S1(i,j1);
5 <some-pragmas-for-loop-j2>
6 for(j2=lbj2(i);j2<=ubj2(i);j2++){S2(i,j2);S3(i,j2);}}
The summary AST for this loop is simply built by creating one node per for loop and one per statement Sx, the body of a loop is made of loops and/or statements, and their nodes are children of said loop in the tree, listed in their syntactic order. For example, above, utilizing a constructor notation, it gives: $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$. Then, a simple rewrites of this tree using loop properties and composition operators will lead to the proposed analytical model, as outlined below.
We now describe the loop properties we associate to each loop. We consider combinations of the following pragmas, based on Merlin’s optimizations, for the loop with iterator i: #pragma ACCEL parallel <factor=ufi>, #pragma ACCEL pipeline <II=IIi>, #pragma ACCEL tile and #pragma ACCEL cache <variable=a>.
We therefore associate to each loop i in the program a property vector that informs about the optimizations to be considered. We define $\vec{P}V_{i}$ as follows: $\vec{P}V_{i}:<ispipelined_{i},II_{i},uf_{i},tile_{i},TC^{min}_{i},TC^{max}_{i}>$ where we have: $ispipelined_{i}=1$ if the loop is pipelined, 0 otherwise; $II_{i}$ is the initiation interval, set to 1 by default; $uf_{i}$ is the parallelism/unroll factor, set to 1 by default (no #pragma ACCEL parallel pragma) and set to $TC^{max}_{i}$ if parallel is defined without a factor $uf_{i}$ specified. $tile_{i}$ is the TC of the innermost loop after strip mining. $TC^{min}_{i}$ is the minimal trip count of loop i, for any of its execution in the program. We also compute the maximal trip count over all executions. These values are computed using polyhedral analysis on the loops (Pouchet et al., 2013a). The pragma cache transfers above the loop i the data needed for the computation of this loop nest for the array a.
This vector is built by syntactic analysis on the program, where the default value $PV_{i}:<0,1,1,1,TC^{min}_{i},TC^{max}_{i}>$ is used for a loop without any pragma. Once all loops have been annotated by their $\vec{P}V$ properties, subsequent treatment can be implemented to mirror the optimizations implemented by the back-end tool.
Modeling Vitis optimizations  AMD/Xilinx Vitis will apply several optimizations automatically, such as auto-pipeline and auto-loop-flatten, some other optimizations when the user gives a compilation option such as tree reduction.
Only a loop with a constant $TC$, i.e., $TC_{max}=TC_{min}$ can be unrolled. The unroll pragma options allow to specify the unrolling factor, $uf_{i}$. When the factor is not specified, it implies that the factor is equal to the $TC$ of the loop. When a loop is pipelined, all innermost loops are automatically fully unrolled. Hence, we also propagate unrolling information, e.g., to mark a full loop nest for full unrolling if an outer loop is marked with #pragma ACCEL pipeline. The pipeline pragma options allow to specify the objective II the user wants to achieve. When II is not specified, it is automatically set to 1. In addition, Vitis will auto-pipeline with a target II of 1 the innermost loop which are not fully unrolled for each nested loop.
Within Vitis, users can enable optimizations like logarithmic time reduction through tree reduction. This optimization choice will be a global option within our model, applicable across the entire model rather than being limited to a specific loop.
Modeling Merlin optimizations  Finally, Merlin will also add automatic optimizations. It will explicitly strip-mine a loop when it is partially unrolled with the innermost loop having a $TC$ equal to the factor and the unrolled pragma applied to that loop. Similarly to Vitis, Merlin will auto-pipeline. Merlin also applies a program transformation for certain pragmas. When there are two perfectly nested and partially unrolled loops, Merlin swaps the two loops strip-mined (if legal), unrolled innermost and flatten and pipeline the two outermost loops. Further, Merlin will automatically transfer the data from off-chip to on-chip and cache on-chip with packing, with a maximum packing of 512 bits for our FPGA while computing if the footprint of the data fit on-chip by static analysis. The pragma tile allows Merlin to strip mine a loop and give the compiler the opportunity to transfer less/more data while respecting resource constraints. For our model we suppose an optimistic data transfer i.e., all memory transfers are done with a packing of 512 bits and each data are transferred once (perfect data reuse).
Consequently, the set of possible $\vec{PV_{l}}$ vectors are adjusted by analyzing the input code, and modifying their initial value, possibly further constraining the set of possible $\vec{PV_{l}}$ based on which program transformation will be performed, as described above. Overall, the $\vec{PV_{l}}$ vectors, along with the summarized AST, contain sufficient information to capture several source-to-source transformations performed by the Merlin compiler for coarse- and fine-grain parallelization, and reason on the likeliness of the optimization to succeed at HLS time (e.g., capturing loops with non-constant trip count).
4. Theoretical Latency and Resource Modeling
We now outline key elements of our analytical performance model, and the associated proofs this model computes a lower bound on latency under resource constraints. Full proofs cover several pages, and are available in the Appendix : we limit below to summarizing the hypothesis and overall proof approach.
4.1. Analytical Model Template
We take as input the summary AST of the program in constructor form, and the set of $\vec{P}V_{i}$ vectors for each loop, as defined in Sec. 3.1. Returning to our example in Sec. 3.1, this summary AST, capturing the loop structure and all statements, is $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$.
Let $I^{\vec{PV_{i}}}_{i}(X)$ be an operator that computes the latency of a region of a code represented by $X$, where $X$ is composed of loops and/or statements. This function is computed as follows: $I^{\vec{PV_{i}}}_{i}(X)=\lfloor II_{i}*(TC^{avg}_{i}/uf_{i}-pip_{i})\rfloor\odot X$ where $\odot$ is $+$ if the loop $i$ is pipelined ($pip_{i}=1$), $\times$ otherwise. Let $C_{i}^{\vec{PV_{i}}}(X_{1},...,X_{n})$ the operators which composed the region, defined as follows: $C_{i}^{\vec{PV_{i}}}(X_{1},...,X_{n})$ is $max_{i=1}^{N}(X_{i})$ if there is no dependency (WaR, RaW, WaW) between the statements of the regions and $\sum_{i=1}^{n}X_{i}$ between the statements which have at least one dependency. Finally, $SL^{\vec{PV_{i}}}_{i}(\vec{S_{k}})$ denotes a region of straight-line code (e.g., an inner-loop body). Intuitively, $SL$ will represent a lower bound on the latency of a code block such that, by the composition across all loops as per the template formula, the result remains a lower bound on the full program latency.
We build the analytical formula template by simple substitution of the operators in the summary AST by the following operators: (1) for every $Loop_{i}(\vec{X})$, replace it by $I^{\vec{P}V_{i}}_{i}(\vec{X})$ ; (2) for every list $\vec{I}$ containing the $I^{\vec{P}V_{i}}_{i}(\vec{X})$ of the step 1, replace it by $C^{\vec{PV_{i}}}_{i}(\vec{I})$; and (3) replace statement lists inside loop i $\vec{S}_{x}$ by $SL^{\vec{PV_{i}}}_{i}(\vec{S}_{x})$. That is for our example, we simply rewrite: $Loop_{i}(Loop_{j1}(S1),Loop_{j2}(S2,S3))$ into the following formula:
$I^{\vec{P}V_{i}}_{i}(C^{\vec{P}V_{i}}(I^{\vec{P}V_{j1}}_{j1}(SL^{\vec{P}V_{j1}}_{j1}(S1)),I^{\vec{P}V_{j2}}_{j2}(SL^{\vec{P}V_{j2}}_{j2}(S2,S3))))$
4.2. Restrictions and Applicability
Our objective is to formulate a lower bound on the latency of a program, after HLS. We therefore include several restrictions:
  • The input program is a pure polyhedral program (Girbal et al., 2006), and its analysis (loop trip counts for every loop, all data dependencies (Feautrier, 1991)) is exact.
  • No HLS optimization shall change the number of operations in the computation: strength reduction, common sub-expression elimination, etc., shall either first be performed in the input program before analysis, or not be performed by the HLS toolchain.
  • The program does not contain ”useless” operations. A typical case is when a value is overwritten before being utilized, e.g., x = 42+51; x = 0;.
  • We only model DSP and BRAM resources for the considered kernel, ignoring all other resources.
  • We assume resources (DSP) sharing across different operations executing at the same cycle is not possible.
4.3. Latency Lower Bound
An important term is $SL$, a latency lower bound for a region of straight-line code. To maintain a lower bound on latency by composition, we operate on a representation of (parts of the) program which is both schedule-independent and storage-independent: the operation graph, or CDAG (Elango et al., 2015). Indeed, a lower bound on this representation is necessarily valid under any schedule and storage eventually implemented and can be used to prove I/O lower bounds on programs (Elango et al., 2015) which is the directed acyclic graph with one node per operation in the code region, connecting all producer and consumer operations to build the operation graph. Then, we can easily compute the length of its critical path, which represents the minimal set of operations to execute serially.
Then, we can build a lower bound on the number of cycles a region $L$ may require to execute, under fixed resources, by simply taking the maximum between the weighted critical path and the work to execute normalized by the resources available.
Theorem 4.1 (Latency Lower Bound under Operation Resource Constraints).
Given $\mathcal{O}$ the set of operation type and $R_{o}$ a count of available resources of type $o\in\mathcal{O}$, $OG$ the operation graph, $cp_{L}$ the critical path of $OG$, and $LO(o)$ the latency function for operation $o$, with $LO(o)\geq 1$. $N_{o}^{L}$ denotes the number of operations of type $o$ in $L$. We define $C_{L}=\sum_{n\in cp_{L}}LO(n)$ the critical path of $L$ weighted by latency of its operations. The minimal latency of a region $L$ is bounded by
$Lat_{R_{o}}^{L}\geq\max(C_{L},\max_{o\in\mathcal{O}}(\lceil N_{o}^{L}\times LO(o)/R_{o}\rceil))$
This theorem provides the building block to our analysis: if reasoning on a straight-line code region, without any loop, then building the operation graph for this region and reasoning on its critical path is sufficient to provide a latency lower bound.
1L1: for (i = 0; i < N; i++) S0: s[i] = 0;
2L2: for (i = 0; i < M; i++) { S1: q[i] = 0;
3 L3: for (j = 0; j < N; j++) {
4 S2: s[j]+=r[i]*A[i][j]; S3: q[i]+=A[i][j]*p[j];}}
Listing 1: Bicg
For example in Lst. 1, if we consider the sub-loop body composed of loops L3 (fully unrolled) and the statements $S2$ and $S3$ as straight-line code regions, we can calculate the critical paths for $S2$ and $S3$ as follows: For $S2$, the critical path is given by: $cp_{S2}=\max(L(+)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$ with $DSP_{o}$ the number of DSP for the operation $o$. For $S3$, the critical path is determined by: $cp_{S3}=\max(L(+)\times\log(N)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$, considering the possibility of a tree reduction. In this context, the critical path for the entire sub-loop body is the maximum of these two individual critical paths, expressed as $\max(cp_{S2},cp_{S3})$.
We now need to integrate loops and enable composition of latency bounds.
Loop Unrolling  Loop unrolling is an HLS optimization that aims to execute multiple iterations of a loop in parallel. Intuitively, for an unroll factor $UF\geq 1$, $UF$ replications of the loop body will be instantiated. If $TC_{l}\mod UF_{l}\neq 0$ then an epilogue code to execute the remaining $TC_{l}\mod UF_{l}$ iterations is needed.
Unrolling can be viewed as a two-step transformation: first strip-mine the loop by the unroll factor, then consider the inner loop obtained to be fully-unrolled. The latency of the resulting sub-program is determined by how the outer-loop generated will be implemented. We assume without additional explicit information this unrolled loop will execute in a non-pipelined, non-parallel fashion. Note this bound requires to build the operation graph for the whole loop body. This is straightforward for inner loops and/or fully unrolled loop nests, but impractical if the loop body contains other loops. We therefore define a weaker, but more practical, bound that enables composition:
Theorem 4.2 (Minimal latency of a partially unrolled loop with factor UF and complex loop bodies).
Given a loop $l$ with trip count $TC_{l}$ and loop body $L$, unroll factor $UF\leq TC$ and $\mathcal{O}$ the set of operation type. Given available resources $R_{o}$ and latencies $L(o)\geq 1$ with $o\in\mathcal{O}$. Then the minimal latency of $l$ if executed in a non-pipelined fashion is bounded by:
$Lat_{R_{o}}^{l,S}\geq\lfloor TC/UF\rfloor*Lat_{R_{o}}^{L}$
The intuition of the proof of Theorem 4.2 is that there cannot exist a shorter critical path in the operation graph of the unrolled loop nest than in the non-unrolled loop body, that is there is no useless operation. This is therefore essential to require the absence of useless operations in the program, as otherwise a shorter critical path may be exposed after loop unrolling, making this theorem incorrect.
Consider the scenario of loop $L0$ within Listing 1, which has been unrolled by a factor denoted as $UF\leq TC_{L0}$ where $TC_{L0}$ is the trip count of $L0$. The latency for one iteration of $S0$ is denoted as $Lat^{S0}_{R_{o}}>0$. In the absence of pipelining, the lower bound of the latency of this sub-loop body is: $\lfloor TC_{L0}/UF\rfloor\times Lat^{S0}_{R_{o}}$.
Loop pipelining  Loop pipelining amounts to overlapping multiple iterations of the loop, so that the next iteration can start prior to the completion of the preceding one. The initiation interval (II) measures in cycles the delay between the start of two consecutive iterations. It is easy to prove our formula template accurately integrates the latency of pipelined loops with the $I$ operator. We compute the minimal II in function of the dependencies of the pipelined loop and the iteration latency of the operations of the statements during the NLP generation. Let $RecMII$ and $ResMII$ be the recurrence constraints and the resource constraints of the pipelined loop, respectively. We have $II\geq\max{(ResMII,RecMII)}$. $RecMII=\max_{i}\lceil\frac{delay(c_{i})}{distance(c_{i})}\rceil$ with $delay(c_{i})$ the total latency in dependency cycle $c_{i}$ and $distance(c_{i})$ the total distance in dependency cycle $c_{i}$. We suppose that $ResMII=1$, as we don’t know how the resource will be used by the compiler. Hence, if the loop is a reduction loop then the $II\geq\frac{IL_{reduction}}{1}$ with $IL_{reduction}$ the iteration latency of the operation of reduction. For a kernel like $for(j=0;j<N;j++)y[j]=y[j-2]+3;$ the $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
Program latency lower bound
We now focus on having the program latency lower bound, with resource constraints. This bound takes into account the limitations imposed by available resources, which can significantly affect the achievable performance. We assume here that the resources consumed are only consumed by the computing units and resource use by the computational unit of one operation can not be reused by the computational unit of another operation executing at the same time. We also assume that the compilers have implemented the pragma configuration given as input.
For DSPs we suppose we have a perfect reuse, i.e., that the computation units for the same operation can be reused as soon as the computation unit is not in use. Under-estimating the resources used is fundamental to proving the latency lower bound, as otherwise another design that consumes less resources than predicted may be feasible, itself possibly leading to a better latency. The lower bound of a kernel is therefore the lower bound of the configuration such that the minimal resource consumption is less than or equal to the resource available.
Theorem 4.3 (Lower bound of the kernel under resource constraint).
Given a loop body $L$, the set of statements $\mathcal{S}$, $\mathcal{O}$ the set of operation type, $N_{o}^{s}$ the number of operations $o\in\mathcal{O}$ for the statements $s$, $DSP_{o}$ the number of resources (DSPs) used for the operation $o\in\mathcal{O}$, $MCU_{o}^{s}$ the maximal computational unit the statement $s$ can use in parallel and the configuration of pragma which is constituted of the vectors $\vec{PV_{i}}$ for each loop. The configuration of pragma, which has the minimal lower bound and which respect $R_{used}^{min}=\sum_{o}\max_{s\in\mathcal{S}}(N_{o}^{s}\times DSP_{o}\times MCU_{o}^{s})\leq R_{o}$, is the lower bound of the kernel.
The intuition of the proof of the Theorem 4.3 is that $R_{used}^{min}$ is the minimal resource that the kernel can use for the configuration of pragma. This implies that if the $R_{used}^{min}>R_{o}$ the configuration over-utilizes resources. Hence, the lowest lower bound which does not over-utilize the resource is the lower bound of the kernel.
5. Non-Linear Formulation for Pragma Insertion
We now present the complete set of constraints and variables employed to encode the latency and resource model as a non-linear program. This section presents a modeling of section 4 in the practical case.
Let $\mathcal{L}$ be the set of loops, $\mathcal{A}$ the set of arrays, $\mathcal{S}$ the set of statements and $\mathcal{O}_{s}$ the operations of the statements $s$. In order to have an accurate model we distinguish for each statement the operation which can be done in parallel, i.e., does not have any loop-carried dependence, $\mathcal{O}_{s_{par}}$ and the reduction operations, i.e., associative/commutative operators to reduce one or more values into a single value, leading to loop-carried dependencies, $\mathcal{O}_{s_{red}}$.
Let $\mathcal{P}$ be the set of different possible pipeline configurations. Let $\forall p\in\mathcal{P}$ define $\mathcal{L}^{p}_{pip}$ the set of loops pipelined and $\forall l\in\mathcal{L}^{p}_{pip}$, $\mathcal{L}^{p}_{under\_pip_{l}}$ the set of loops under a loop pipelined and $\mathcal{L}^{p}_{above\_pip_{l}}$ the set of loops above the loop pipelined $l$. Let $\forall s\in\mathcal{S}$ define the set of nested loops which iterate the statement $s$, $\mathcal{L}_{s}$. $\forall a\in\mathcal{A}$ and for $d$ a dimension of the array $a$, let $\mathcal{C}_{a_{d}}$ be the set of loops which iterates the array $a$ at the dimension $d$. $\forall l\in\mathcal{L}$, let designate $d_{l}$ the maximum dependency distance of the loop $l$. And let $II_{s}$ be the II of the loop pipelined for the statement $s$.
The II for each loop, the dependencies, the properties of the loops, the TC, the iteration latency of the parallel operations and the reduction operations and the number of DSPs per operation per statements are computed at compile time with PolyOpt-HLS (Pouchet et al., 2013a) and used as constants in the NLP problem.
5.1. Variables
Variables in the formulation correspond to $PV_{l}$ attributes. We consider the possibilities of pipelining (Eq. 3), unrolling (Eq. 1), and tiling (Eq. 2) for each loop. Additionally, we include the possibility of caching an array that is iterated over by the loop (Eq. 4).
5.2. Modeling Compiler Transformations
Now that we have defined our design space, we need to constrain the space by removing infeasible cases and those that do not comply with the rules of the compilers.
Pipeline Rules  Vitis HLS unrolls all loops under the pipelined loop. This implies that all loop $l$ under the pipelined loop must have $loop_{l}\_UF==TC_{l}$ (Eq. 15). Considering this constraint, it is important to note that only one pipelined loop can exist per nested loop (Eq. 5). If multiple pipelined loops were present, the loops beneath the first pipelined loop would be unrolled instead.
Memory Transfer Rules  Merlin automates the process of transferring data on-chip and applying array partitioning. The tool caches data on-chip and packs it in chunks of up to 512 bits, enabling efficient transfer speed. When the data is already present on-chip it can be reused provided that resource constraints are satisfied. The compiler caches on-chip the data only above the loop pipelined (Eq. 14).
Dependencies  Loop-carried dependencies are managed using constraints (Eq 8). If a loop has a dependency distance of $n$, this means that if we unroll with an unroll factor $uf>n$ this is equivalent to unrolling the loop with a factor $uf=n$ because the statements corresponding to the iteration $\{uf+1,\dots,n\}$ will be executed only after the first n statements are executed due to the dependency. Loop-independent data dependencies are managed at the objective function, as elaborated in Section 5.4.
Supplementary Rules  In addition, we add the constraints for the maximum unrolling (Eq. 10), the divisibility of the problem size of the unroll factors (Eq. 6), the tile size (Eq. 7) and the maximum array partitioning (Eq. 13). For this last, this corresponds to adding an upper bound to the product of the UF of all the loops which iterate the same array on different dimensions.
During our DSE, we can force the solution to be fine-grained. In this case we add a constraint where the loop above the loop pipelined has a UF of 1, i.e., for all loop $l$ above the pipelined loop $loop_{l}\_UF==1$ (Eq. 9). We also constrain the resources, modeling their sharing optimistically. We consider the number of DSPs (Eq. 11) and on-chip memory (Eq. 12) used. As the consumption of DSPs can be difficult to estimate due to resource sharing we utilize an optimistic estimate, which considers a perfect reuse/sharing: as soon as a computation unit is free, its resource can be reused.
5.3. Constraints
(1) $\forall l\in\mathcal{L},1\leq loop_{l}\_UF\leq TC_{l}$
(2) $\forall l\in\mathcal{L},1\leq loop_{l}\_tile\leq TC_{l}$
(3) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
(4) $\forall l\in\mathcal{L},\forall a\in\mathcal{A},loop_{l}\_cache\_array_{a}\in\{0,1\}$
(5) $\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\leq 1$
(6) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l}==0$
(7) $\forall l\in\mathcal{L},loop_{l}\_tile\%TC_{l}==0$
(8) $\forall l\in\mathcal{L},\text{ if }dd_{l}>1,loop_{l}\_UF\leq d_{l}$
(9) $\forall l\in\mathcal{L},\forall l^{\prime}\in\mathcal{L}_{above\_l},loop_{l}\_pip*loop_{l^{\prime}}\_UF<=1$
(10) $\forall s\in\mathcal{S},\prod_{l\in{L}_{s}}loop_{l}\_UF\leq MAX\_PARTITIONING$
(11) $DSPs\_used_{optimistic}=\sum_{op\in\{+,-,*,/\}}\max_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
(12) $\begin{cases}&\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}loop_{l}\_cache\_array_{a}\\ &\times footprint\_array_{a}\_loop_{l}\leq Mem\end{cases}$
(13) $\begin{cases}&\forall a\in\mathcal{A},\forall(d,d^{\prime})\in\mathbb{N}^{2}\text{ with }d\neq d^{\prime},\forall l\in C_{a_{d}},\forall l^{\prime}\in C_{a_{d^{\prime}}},\\ &loop_{l}\_UF\times loop_{l^{\prime}}\_UF\leq MAX\_PARTITIONING\end{cases}$
(14) $\begin{cases}&\forall p\in\mathcal{P},\forall lp\in\mathcal{L}^{p}_{pip},\forall lbp\in\mathcal{L}_{under\_pip_{lp}}\forall a\in\mathcal{A},\\ &loop_{lbp}\_cache\_array_{a}==0\end{cases}$
(15) $\begin{cases}&\forall p\in\mathcal{P},\forall lp\in\mathcal{L}^{p}_{pip},\forall lbp\in\mathcal{L}_{under\_pip_{lp}},\\ &loop_{lp}\times loop_{lbp}\_UF==loop_{lp}\times TC_{lbp}\\ \end{cases}$
5.4. Objective function
Lastly, we need to define the objective function ($obj\_func$) that supports fine-grained and coarse-grained parallelism. Fine-grained parallelism involves duplicating a specific statement(s), while coarse-grained parallelism duplicates modules, including statements and loops. However, it may not always be feasible to achieve parallelism based on the characteristics of the loops and the level of parallelism required. Therefore, we distinguish between parallel and reduction loops. A parallel loop can be coarse and fine-grained unrolled, whereas a reduction loop can only be fine-grained unrolled with a tree reduction process that operates in logarithmic time.
As the pragmas cache are part of the space we compute the communication latency with these pragmas. If more than one array is transferred above the same loop we take the maximum as Merlin transferred them in parallel. To ensure these properties, we formulate the objective function for each pipeline configuration. The objective function uses the combined latencies of communication and computation. When using Merlin, communication and computation do not overlap, but communication tasks can overlap when they occur consecutively in the code at the same level. Consequently, for each loop, where two arrays are transferred consecutively within the loop, we calculate the sum of the maximum latencies for these transferred arrays ($L_{mem}$).
In every loop nest, there will invariably be a pipeline loop due to either user-inserted or compiler-inserted instructions (AMD/Xilinx Merlin and Vitis automatically insert the pragma pipeline if it is not done by the user or the previous compiler). Therefore, the objective function takes the following form: $TC_{ap}\times(IL+II\times(\frac{TC}{UF}-1))$, where $TC_{ap}$ includes the loops situated above the pipeline. Parallel loops above the pipeline can be coarse-grained parallelized. The iteration latency within the unrolled loop body is divided into either reduction operations or non-reduction operations, as reduction operations require logarithmic time for the reduction process. The variable IL encompasses the latencies of the statements found within the pipelined loop body. Independent statements can be executed in parallel and statements with dependencies are summed, as detailed in Section 4.1.
$\left.\begin{cases}&TC_{ap}=\prod_{l\in\mathcal{L}^{par}_{above\_pip}}\frac{TC_{l}}{loop_{l}\_UF}\times\prod_{l\in\mathcal{L}^{red}_{above\_pip}}TC_{l}\par\\ &IL=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}_{under\_pip}}\frac{TC_{l}}{loop_{l}\_UF}\times\log_{2}(loop_{l}\_UF)\\ &L_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(loop_{l}\_cache_{a}\times footprint_{a}\_loop_{l})\\ &obj\_func=TC_{ap}\times(IL+II\times(\frac{TC_{lp}}{loop_{lp}\_UF}-1))+L_{mem}\end{cases}\right.$
5.5. Example
1Loop0: for(i=0; i<2100; i++) S0:y[i] = 0;
2Loop1: for(i=0; i<1900; i++) { S1:t[i] = 0;
3 Loop2: for(j=0; j<2100; j++) S2:t[i]+=A[i][j]*x[j];
4 Loop3: for(j=0; j<2100; j++) S3:y[j]+=A[i][j]*t[i];}
Listing 2: Atax Large
We now employ the Atax kernel as an illustration. $S0$ and $S3$ do not have inter-iteration dependencies within their respective loops, namely $Loop0$ and $Loop3$. Therefore, it is possible to pipeline $Loop0$ and $Loop3$ with an $II\geq 1$. In other words, for all iterations within these loops, there are no dependencies on previous iterations within the same loop. Loop1 and Loop2 are reduction loops, and the reduction operations is an addition which has an IL of $IL_{+}$ cycles. So the II ¿= $IL_{+}$. If loop1 is pipelined, there is a dependency between S1, S2 and S3 Between S1 and S2 the distance is 1, so we just add IL_S1 and IL_S2. Between S2 and S3 the distance is log(N) so the final equation will be IL_S1 + IL_S2 * log(N) + IL_S3 for the loop body cycle. If statements can be run at the same time (i.e., there is no dependency) it is a max instead of an addition. If we encounter code like $for(j=0;j<N;j++)y[j]=y[j-2]+3;$, a straightforward approach to handling this type of dependency is to impose a constraint such as $loop_{l}\_UF\leq 2$, which is represented by equation 8. In this case, due to dependencies an $UF>2$ is similar to $UF=2$.
6. Design Space Exploration
We now present our Design Space Exploration (DSE) approach. Our approach focuses on identifying designs with the most promising theoretical latency within the available design space. However, it may result in suboptimal designs if the selected pragmas are not applied during compilation. To address this potential issue and ensure high QoR, we conduct an additional exploration within a restricted subspace. Our DSE explores two additional parameters: the type of parallelism and the maximum array partitioning factor. Array partitioning is a technique commonly used in FPGA contexts to divide arrays or matrices into smaller sub-arrays, which can be stored in independent memory blocks known as Block RAMs (BRAMs). AMD/Xilinx HLS has a limit of 1,024 partitions per array. The array partitioning is calculated by taking the product of loops that iterate the same arrays on different dimensions (cf. Section 5). So constraining the maximal array partitioning also constrains the maximal UF. This NLP based DSE technique is presented in the Algorithm 1. The DSE starts without constraint on parallelism and array partitioning. Then we alternate constraints on parallelism while decreasing the maximum unrolling factor and array partitioning.
Data: $kernel$ // without Pragma
Data: $timeout\_HLS$, $timeout\_NLP$
Result: $kernel$ // with Merlin Pragma
$nlp\_file\leftarrow generate\_nlp\_file(kernel)$ , $min\_lat\leftarrow\infty$;
for $max\_array\_partitioning\in\{\infty,2048,1024,512,256,128,64,32,16,8,1\}$ do
       for $parallelism\in\{coarse+fine,fine\}$ do
             $current\_nlp\_file\leftarrow$ change_max_array_partitioning($copy(nlp\_file)$, $max\_array\_partitioning$);
             if $parallelism==fine$ then
                   add_constraint_only_fine_grained_parallelism( $current\_nlp\_file$);
             end if
            $pragma\_configuration,lower\_bound\leftarrow SOLVER(current\_nlp\_file,timeout\_NLP)$ ;
             if $lower\_bound<min\_lat$ then
                   $current\_kernel\leftarrow$ introduce_pragma($copy(kernel)$, $pragma\_configuration$);
                   $hls\_lat,valid\leftarrow MERLIN(kernel,timeout\_HLS)$ ;
                   if $valid$ then // no over-utilization
                        $min\_lat\leftarrow min(min\_lat,hls\_lat)$;
                   end if
             end if
       end for
end for
Algorithm 1 NLP-DSE
In order to reduce the maximum unroll factor and array partitioning, we modify the parameters specified in the NLP file. And we automatically add constraints (Eq 9) to restrict parallelism to fine-grained levels, as described in Section 5. The choice to restrict the maximum array partitioning to the power of 2 is to improve the speed of the DSE. Adding possibilities would permit exploring a larger space and potentially finding a design with a faster latency at the cost of a longer DSE.
7. Evaluation
We now present our experimental results using a set of polyhedral computation kernels.
7.1. Setup
We use kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]). In addition we add a kernel of Convolution Neural Network (CNN). A single-precision floating point is utilized as the default data type in computations. Computations operate on medium and large datasets from PolyBench/C (PolyBench, [n. d.]) in order to have kernels with large footprint and have a large enough space to explore. The problem size and loop order of CNN are J,I=256, P,Q=5 H,W=224. A description of each benchmark can be found in Table 2. The ludcmp, deriche and nusinnov kernels are not present as PolyOpt-HLS (Pouchet et al., 2013b) does not handle negative loop stride. Cholesky and correlation contains a sqrt() operation which we do not support currently. Finally, we removed FDTD-2D because it exposed a bug in Merlin, and this generated a program where data dependencies are not fully preserved.
We evaluate designs with AMD/Xilinx Merlin (Xilinx, 2023a). The synthesis is carried out with AMD/Xilinx Vitis 2021.1. We choose the option ”-funsafe-math-optimizations” to enable commutative/associative reduction operators and implementation of reductions in logarithmic time. We change the default on-chip memory size of Merlin by the size of the device we use. As the target hardware platform, we run the Xilinx Alveo U200 device where the target frequency is 250 MHz.
We analyze the kernels and automatically generate each NLP problem with a version of PolyOpt-HLS (Pouchet et al., 2013a). We modified and extended for our work. Employing the AMPL description language to solve the NLP problems, we ran the commercial BARON solver (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) version 21.1.13. For our experiments, we utilize 2 Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz and 252GB DDR4 memory.
7.2. Experimental Evaluation
We compare our method with AutoDSE (Sohrabizadeh et al., 2021), described in Section 2, and we automatically generate the space of AutoDSE with the command ds_generator. We replace the UF and tile size by all the UF and tile size which divide the TC in order to have the same space. AutoDSE doesn’t impose any constraints on parallelism or the maximum array partitioning. It employs an incremental exploration approach, enabling it to make compiler-specific pragma selections.
Table 2 displays the space size of each design. The DSE is done in 4 parts with two threads for each (default parameter), with a timeout for the generation of the HLS report of 180 minutes, and a timeout of the DSE of 600 minutes (not always respected cf. Table 2). For our method we take the same parameters, and we add a timeout for BARON of 30 minutes.
7.3. Comparison with AutoDSE
FS NLP-DSE AutoDSE Perf. Improvement
Kernel NL ND S Space S GF/s GF/s T DE DT GF/s T DE DT ER T GF/s
covariance 7 34 M 1.80E+11 0.08 0.75 336 21 8 0.28 645 161 15 115 1.92x 2.64x
covariance 7 34 L 1.92E+13 0.39 0.73 466 21 9 0.62 2,849 209 68 118 6.11x 1.16x
2mm 6 13 M 1.37E+10 13.19 117.48 70 18 0 0.41 1,870 101 37 49 26.71x 288.88x
2mm 6 13 L 1.15E+12 0.57 2.17 456 17 3 0.40 1,835 291 38 240 4.02x 5.41x
3mm 9 19 M 1.20E+15 13.86 138.73 242 18 0 0.39 698 82 15 57 2.88x 354.12x
3mm 9 19 L 6.18E+17 0.18 1.01 486 19 3 0.59 968 112 18 81 1.99x 1.71x
atax 4 12 M 1.40E+05 1.96 1.96 194 10 1 1.98 1,653 175 13 136 8.52x 0.99x
atax 4 12 L 1.60E+07 0.47 1.52 205 11 2 0.44 2,325 166 30 106 11.34x 3.46x
bicg 3 10 M 1.90E+04 0.99 0.99 248 12 1 0.98 729 65 2 28 2.94x 1.01x
bicg 3 10 L 4.44E+05 1.68 1.68 218 12 1 0.50 3,355 236 42 176 15.39x 3.38x
cnn 6 2 - 6.43E+06 0.39 97.99 213 16 1 97.99 1,292 28 19 480 6.06x 1.00x
doitgen 5 30 M 8.64E+06 19.75 19.75 193 13 0 18.95 819 296 14 248 4.24x 1.04x
doitgen 5 30 L 3.63E+07 0.08 102.62 241 20 1 110.66 1,299 222 24 169 5.39x 0.93x
durbin 4 55 M 1.08E+02 0.01 0.20 193 7 4 0.12 134 25 0 23 0.69x 1.65x
durbin 4 55 L 9.00E+00 0.12 0.12 212 3 1 0.12 31 7 0 5 0.15x 1.00x
gemm 4 6 M 2.30E+06 105.18 105.18 185 21 1 68.91 1,345 86 27 34 7.27x 1.53x
gemm 4 6 L 1.47E+07 32.98 32.98 450 18 7 2.77 2,810 188 47 133 6.24x 11.89x
gemver 7 13 M 7.72E+11 0.78 9.45 218 21 4 2.99 847 65 5 28 3.89x 3.16x
gemver 7 13 L 1.28E+13 9.94 9.94 290 21 7 0.18 1,756 221 206 10 6.06x 54.69x
gesummv 2 17 M 6.12E+02 1.97 1.97 220 14 3 1.97 836 80 8 47 3.80x 1.00x
gesummv 2 17 L 6.33E+03 1.82 2.64 236 18 1 0.56 692 94 29 60 2.93x 4.73x
gramschmidt 6 34 M 1.75E+07 1.58 1.58 364 7 3 0.44 934 109 92 8 2.57x 3.56x
gramschmidt 6 34 L 1.22E+08 2.34 2.34 420 6 4 0.95 819 265 11 239 1.95x 2.47x
lu 5 16 M 2.28E+03 0.03 0.03 614 19 11 0.04 849 193 1 159 1.38x 0.98x
lu 5 16 L 3.99E+03 0.03 0.04 335 9 2 0.03 812 258 5 219 2.42x 1.03x
mvt 4 6 M 1.38E+07 7.77 7.77 212 17 1 7.77 893 166 6 106 4.21x 1.00x
mvt 4 6 L 7.41E+07 12.90 12.90 181 20 3 1.10 1,240 249 20 189 6.85x 11.78x
symm 3 33 M 2.31E+04 0.04 0.20 63 5 0 0.20 691 142 35 89 10.97x 1.00x
symm 3 33 L 6.64E+04 0.21 0.31 540 8 5 0.21 612 731 0 692 1.13x 1.52x
syr2k 4 6 M 2.32E+04 0.07 1.74 224 16 2 1.20 685 230 17 203 3.06x 1.45x
syr2k 4 6 L 6.67E+04 1.30 1.42 420 15 7 1.30 768 293 21 262 1.83x 1.09x
syrk 4 6 M 2.32E+04 0.49 1.32 224 16 2 0.61 631 280 4 264 2.82x 2.15x
syrk 4 6 L 6.67E+04 0.94 2.07 466 17 7 0.65 643 410 0 398 1.38x 3.16x
trisolv 2 13 M 3.60E+02 0.03 0.03 69 12 0 0.04 694 69 33 23 10.06x 0.98x
trisolv 2 13 L 6.30E+02 0.04 0.04 75 18 0 0.04 651 127 2 98 8.68x 0.99x
trmm 3 8 M 2.31E+04 0.01 0.05 20 16 0 0.04 630 401 4 367 31.50x 1.29x
trmm 3 8 L 6.64E+04 0.02 0.06 425 17 2 0.03 760 167 159 4 1.79x 1.79x
floyd-warshall 3 21 M 8.65E+04 0.17 0.61 246 17 3 0.10 1,605 60 22 29 6.52x 6.15x
floyd-warshall 3 21 L 3.29E+06 0.17 1.31 381 20 6 0.10 2,728 150 71 77 7.16x 13.09x
heat-3d 7 42 M 3.04E+07 0.23 3.75 402 17 7 3.75 928 75 33 35 2.31x 1.00x
heat-3d 7 42 L 4.37E+07 0.13 0.62 520 13 6 0.63 740 109 70 35 1.42x 0.99x
jacobi-1d 3 14 M 1.48E+03 11.43 11.43 55 4 0 11.53 948 126 3 74 17.24x 0.99x
jacobi-1d 3 14 L 4.00E+04 5.95 5.95 386 9 6 2.65 1,283 173 21 123 3.32x 2.25x
jacobi-2d 5 22 M 8.07E+06 0.20 3.32 379 12 5 3.32 674 158 26 98 1.78x 1.00x
jacobi-2d 5 22 L 1.14E+07 0.26 1.25 427 9 6 1.25 1,106 231 39 171 2.59x 1.00x
seidel-2d 3 27 M 4.26E+03 0.05 0.05 365 5 1 0.05 796 103 27 68 2.18x 1.01x
seidel-2d 3 27 L 1.77E+05 0.05 0.05 540 13 7 0.05 880 91 33 48 1.63x 1.00x
Average 5.38 15.11 296 7.44 1,123 5.69x 17.24x
Geo. Mean 0.51 1.54 247 0.65 916 3.70x 2.38x
Table 2. Comparison with AutoDSE
Table 2 shows the comparison with AutoDSE. NL, ND, S, and Space S are respectively the number of loops, the number of polyhedral dependencies (WaR, WaW, RaW), problem size (L for Large and M for Medium) and space size. For each method we compute the throughput (GF/s) in GFLOPs per second, the total time of the DSE (T) in minutes, the number of designs explored (DE) and the number of designs timeout (DT). In addition, for AutoDSE we add the number of design that are early rejected/prune (ER) as AutoDSE prunes the design when AMD/Xilinx Merlin can not apply one of the pragmas, due to its analysis limitations.
To illustrate the performance achievable without a complete DSE, the first synthesizable design found with NLP-DSE (FS) is displayed. Indeed, due to our under-estimation of resources, the theoretically best design produced by NLP solving may not be synthesizable. We report the improvements in DSE time (T) and throughput (GF/s).
The performance of the kernel evaluated show significant improvements in both time and throughput. The time of the DSE is 5.69x faster on average (3.70x for geo-mean) and the throughput is 17.24x higher on average (2.38x for geo-mean) for the kernel evaluated. For almost all (46/47) kernels and problem sizes the method identifies a design with a throughput similar to (+/- 2%), or better than, AutoDSE. We have a slight slowdown for Doitgen Large because NLP-DSE explores the design found by the NLP with a maximum array partitioning of 2048 which timeouts, and then 1024 which is the best design found. However AutoDSE finds a design with a maximum array partitioning of 1280. By changing the maximum array partitioning to 1280 we find the same configuration as AutoDSE. Thus it is possible to obtain designs with a better performance but at the cost of a longer search. For all kernels and problem sizes, except Durbin, NLP-DSE is faster than AutoDSE. AutoDSE prunes all configurations of Durbin which explains the speed of AutoDSE for this kernel.
We can observe a difference of the performance for the same kernel in function of the problem size. If we take the example of 2mm, the difference has many factors. First as the footprint of the kernel becomes more important, it begins overusing the BRAMs. A large parallelism requires a bigger array partitioning which considerably increases the number of BRAMs and uses more BRAMs than available. Additionally, for large problems with high levels of parallelism, there are multiple instances of timeouts observed. Furthermore, the compilers applied the pragmas more efficiently for smaller problem sizes. We observe twice as many kernels where the pragmas are not applied as expected for the large problem size.
For Atax Large (Lst. 2), AutoDSE explores 166 designs of which 106 are early rejected and 30 timeout. AutoDSE starts by partially unrolling Loops 2 and 3 and will then attempt to do a coarse-grained parallelization on Loop 1 with all divisors, which is impossible due to dependencies. Although AutoDSE manages to prune/early reject the designs because Merlin cannot apply the pragmas, it still requires several minutes of compilation by Merlin for each unroll factor. In parallel, AutoDSE tries to pipeline the outermost loops (and therefore unroll the innermost loops) which creates numerous timeouts. Although the first two designs timeout due to too high level of parallelism, NLP-DSE allows us to find a configuration with the innermost loops unrolled with a UF=700. This allows us to find a design with a 3.46x higher throughput in 11.34x less time.
Our method experiences some timeouts for designs with high levels of parallelism. However, thanks to our DSE approach, we quickly identify optimized designs where each loop body has a similar level of parallelism. For 20/47 cases, the first synthesizable design is equal to the best design of the DSE. This is because compilers can be conservative and not apply pragmas as expected. In this case, another configuration is applied than what was identified by the NLP, which explains the difference in performance.
7.4. Accuracy
The tightness of the lower bound estimation relies on the correct application of pragma directives such as pipeline and parallel. It also assumes that Merlin can efficiently transfer memory from off-chip to on-chip using 512-bit chunks. Finally, it assumes that Merlin optimally handles the transfer of memory from off-chip to on-chip. Figures 1a and 1b compare the measured HLS latency for every synthesizable design explored during our DSE with its predicted latency per solving the NLP. The Y-axis represents log(latency) and the X-axis the rank of the design sorted by HLS latency. For Fig 1b we exclude designs when we detect that the pragma parallel and pipeline are not applied as defined by Merlin. We observe that about half of the designs have at least one pragma not applied, leading logically to a larger difference between measured and predicted latency. Generally speaking, for parallelization pragmas, Merlin is more restrictive for coarse-grained parallelization, in many cases these pragmas are not applied. We also observe certain cases where the partitioning is not done correctly which does not allow a pipeline with II=1 when it is theoretically possible. For Figure 1b we observe a better overall accuracy, albeit imperfect. These differences are due in large part to how Merlin eventually implemented memory transfers, which we model optimistically.
(a) when some pragmas are not applied
(b) when all pragmas are applied
Figure 1. Comparison of the HLS Latency vs. Lower Bound
We observe in Fig. 1a and Fig. 1b one configuration where the lower bound property is not maintained (shown in red). This corresponds to a configuration of the Heat-3d kernel, where the pragma loop_flatten has been applied automatically, which changes the program structure. Overall unless Vitis applies loop_flatten automatically, which we do not model, our estimate is a lower bound for the cases evaluated.
7.5. Scalability
To mitigate prolonged solving times for specific kernels and problem sizes, we’ve implemented a 30-minute timeout constraint for the BARON solver. While this timeout doesn’t guarantee achieving optimality, it ensures that the solver provides the best solution it has found within the time limit. In Table 3, we present statistics regarding the number of problem timeouts (ND T/O) and problem non-timeouts (ND NT/O), along with the average time in seconds (Avg Time) for all problems and exclusively for those that did not time out. We can note that the 20 NLP problems for CNN finish in few seconds with an average of 3.71 seconds.
We notice that 12 kernels exhibit at least one NLP problem that times out. To investigate scalability further, we conducted restarts for NLP problems that timed out at 30 minutes, extending the timeout to 30 hours. For 30 out of 126 problems (23.8%), we found an optimal theoretical solution within an average time of 3.13 hours. We observe that problems timing out after 30 hours often involve trip counts with numerous divisors, significantly expanding the space for the unroll factor. Consequently, non-linear conditions involving more than three unknown variables of unroll factors (UFs) become extremely time-consuming to resolve. By relaxing these constraints, we are able to find a solution in seconds but this can result in infeasible designs due to over-utilization of resources as these constraints are removed. For 23.8% of problems not timing out at 30 hours, we examined the disparity in objective function values provided by the solver when it times out at 30 minutes (representing the best solution found so far) versus when it discovers the optimal solution. For 25 out of 30 problems, the estimated latency is exactly the same. However, for the remaining 5 problems, the differences in the estimated latencies range from a mere 0.04% up to 2,426%.
Size ND T/O ND NT/O Avg Time Avg Time NT/O
Medium 7 469 55s 29s
Large 119 361 479s 43s
All 126 830 268s 35s
Table 3. Study of the scalability
8. Related Work
NLP-DSE makes it possible to automatically introduce pragmas in order to obtain a design with a good QoR. Many previous works using different DSE methods have the same objective. Model-free DSEs (Sohrabizadeh et al., 2021; Yu et al., 2018; Ferretti et al., 2018) evaluate each candidate by generating the HLS report. The synthesis time or report generation can last several hours, which considerably reduces the explored design space. Additionally, some DSE, such as (Sohrabizadeh et al., 2021), may lead to local minima. In order to avoid the synthesis time limitation, model-based DSEs and AI-driven DSEs have been developed. They estimate the performance of each design using a cost model (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Wang et al., 2017, 2017; Zhong et al., 2017; Zuo et al., 2015) , Graph Neural Networks (GNN) (Sohrabizadeh et al., 2022; Wu et al., 2023) or decision trees (O’Neal et al., 2018; Yu et al., 2021; Makrani et al., 2019). These DSEs estimate the QoR of the design in a few milliseconds. However, the evaluation of a large number of designs can still take several hours and the accuracy may not match the HLS report. Hence, evaluating only the top-$n$ results with HLS may lead to sub-optimal solutions. Other works allow one-shot optimization with code transformations, e.g., (Ye et al., 2022) but the space of hardware directives and code transformations is limited. Some focus on a predefined micro-architectures, e.g., (Wang et al., 2021) but are limited in their application. NLP-DSE is a hybrid approach: our cost model in the form of NLP makes it possible to consider very large design spaces in a few minutes. This makes it potentially even faster than existing cost models. Yet we rely on HLS to ensure accurate performance.
Recent advances in optimization solvers such as BARON (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) have allowed NLP-based approaches to become a promising alternative to approximate ILP-based methods, as they can encode more complex and realistic performance models. Previous works (Bondhugula et al., 2008; Pouchet et al., 2013a; Zuo et al., 2017) encode a cost model as Linear Programming (LP) problem. However, it requires an approximation, such as approximating the volume of communication inter loops (Bondhugula et al., 2008; Pouchet et al., 2013a), or simplifying the space by exposing direct parallelization in the problem (Zuo et al., 2017).
The selection of tile sizes remains fundamental for the final QoR Similar to our approach, (Liu et al., 2017) uses a cost model to select the tile size. Although their space is much more complete than ours as we are restricted to Merlin’s transformations, their method does not allow the evaluation of the whole space.
9. Conclusion
Our work targets the automatic selection of pragma configurations for HLS with a framework that automatically inserts Merlin pragmas into loop-based programs. Our framework is guided by an analytical performance and resource model, which serves as a lower bound estimation for the achievable performance across all possible configurations.